# references.bib

### v1

@@ Abstract
This paper considers the problem of model selection under domain shift. In this setting, it is proposed that a high maximum mean discrepancy (MMD) between the training and validation sets increases the generalisability of selected models. A data splitting algorithm based on kernel k-means clustering, which maximises this objective, is presented. The algorithm leverages linear programming to control the size, label, and (optionally) group distributions of the splits, and comes with convergence guarantees. The technique consistently outperforms alternative splitting strategies across a range of datasets and training algorithms, for both domain generalisation (DG) and unsupervised domain adaptation (UDA) tasks. Analysis also shows the MMD between the training and validation sets to be strongly rank-correlated ($\rho=0.63$) with test domain accuracy, further substantiating the validity of this approach.
@@ 1 Introduction
The ability for models to maintain high performance on data lying outside their training distribution, known as domain generalisation (DG), is crucial to the widespread deployment of AI. Although extensive research has been conducted towards developing more generalisable training algorithms [Gulrajani2021InGeneralization], significantly less focus has been given to increasing the robustness of the model selection process, despite being as integral a part of the learning problem, and indeed, just as susceptible to distribution shifts, as the fitting of the model itself.
As with model parameters, hyperparameter choices based on in-distribution (ID) performance may not be optimal on an out-of-distribution (OOD) test domain. As such, OOD validation sets (distinct from both the training and test sets) are commonly employed, normally created through the use of additional metadata (e.g., domain labels). Empirical evidence supports the assumption that this encourages the selection of more generalisable hyperparameters [Koh2021WILDS:Shifts]. In this paper, it is proposed that the validation set should indeed be maximally domain shifted from the training set, given some quantitative distance measure such as the maximum mean discrepancy (MMD), while still retaining a relevant distribution. It is proposed that this would further encourage the selection of robust hyperparameters.
@@
To this end, it is noted that the kernel k-means clustering problem can be re-posed as the problem of maximising the empirical MMD between clusters (weighted by cluster size). In light of this, it is proposed to perform a validation split based on kernel k-means, and a modified clustering algorithm is presented for this purpose. Specifically, constraints are introduced to the cluster assignment step to control the holdout fraction (i.e., the cluster sizes), and to preserve class and (optionally) domain/group distributions; this is then formulated and solved as a linear program, providing convergence guarantees not present in prior work. In addition, the Nyström method for low-rank approximation is employed to make the algorithm tractable for large datasets.
@@
The proposed method provides a model selection strategy based on OOD performance that does not require additional metadata, which is not always available. It is also argued that the method is able to capture more nuances in real-world data than can be described by a single domain variable. For example, in tumour identification, domain shifts can be caused by variations in sample preparation methods or patient populations. Or, in wildlife monitoring (both acoustic and visual), these can be due to differences in environmental or weather conditions, data collection equipment, or new unseen events [Koh2021WILDS:Shifts]. However, all this information is unlikely to be described in the available meta-data, which may only state the hospital of origin or the recording location, respectively. In such cases, splitting the data along lines closer to the underlying cause of the shift, as determined by the clustering algorithm, rather than a more loosely correlated proxy, may permit a more informed model selection which can result in hyperparameters better suited to the nature of the shift.
Clustering the development data can also serve as a natural way of disentangling spurious cues and ensuring these are not carried over to the validation set. To illustrate this, consider the Coloured MNIST (CMNIST) dataset [Arjovsky2019InvariantMinimization], in which a digit’s colour is spuriously correlated with its label. Applying vanilla k-means here would perfectly separate the two colour groups. The colour shortcut would then be absent from the validation set, meaning the model selection can be based on recognition of the intended attribute (the digit) instead. Of course, due to CMNIST’s unrealistically simple structure, it is trivial for vanilla clustering to separate the spurious attribute here. Moreover, the validation set would then closely align with the test distribution in this case, which is unlikely to occur in real-world data. Therefore, this dataset, along with any others comprising synthetically varying subpopulations of a spurious variable (CelebA [Sagawa2019DistributionallyGeneralization, Liu2015DeepWild], Waterbirds [Sagawa2019DistributionallyGeneralization], Spawrious [Lynch2023Spawrious:Biases], etc.), are not considered valid benchmarks to test this method on.
@@
  • Description of a constrained clustering algorithm based on kernel k-means which can be used to perform a validation split in applications expected to involve domain shifts.
  • Comparison of this approach with existing validation strategies for a range of datasets and algorithms, in both DG and UDA settings.
@@ 1.1 Relation to existing work
Gulrajani2021InGeneralization reviewed 3 criteria for model selection in a DG setting: ID accuracy (on a randomly held-out subset of each training domain); OOD accuracy on an additional domain held-out using metadata (referred to as leave-one-domain-out); and test-distribution accuracy on a held-out subset of the test domain (referred to as the oracle criterion), which can be used to provide an upper bound on performance. Where the validation set comprises multiple domains, these authors take the average accuracy over these domains, although it has been argued that considering only the worst-performing domain provides a more robust indicator of generalisability [Sagawa2019DistributionallyGeneralization, Gao2023Out-of-DistributionAugmentations].
In a link to the latter, clustering-based splits can also be viewed as a measure of worst-case performance – where the worst-case performance of a given set of hyperparameters is assumed to correspond to the validation accuracy when the MMD between the training and validation sets is maximised.
@@
In addition to validation accuracy, it has been suggested that a model’s stability to distribution shifts should also be explicitly considered. Prior work has quantified stability in terms of the expected calibration error (the average deviation between accuracy and confidence) [Wald2021OnGeneralization]; the MMD between features from different domains [Lyu2023AGeneralization]; or the average variation of each feature between domains [Ye2021TowardsGeneralization].
[…]
1.2 Preliminaries and notation
@@
Let $S=\{(x_{1},y_{1},d_{1}),\ldots,(x_{n},y_{n},d_{n})\}$ be the development set consisting of input-label-domain triplets over $\mathcal{X}\times\mathcal{Y}\times\mathcal{D}^{S}$. Similarly, let $E$ be the evaluation set over $\mathcal{X}\times\mathcal{Y}\times\mathcal{D}^{E}$, where $\mathcal{D}^{S}\cap\mathcal{D}^{E}=\emptyset$ (i.e., the development and evaluation domains are disjoint). For ease of notation, subscripts are used on sets to simplify set-builder notation, in two ways. Firstly, “slices” of a set are denoted using capitals, for example:
@@
It may also be necessary or desirable to control domain distributions – for example, certain training algorithms may require the domains in $T$ to be uniformly represented to avoid overfitting. In this case, the constraints should be taken over all $(y,d)$ pairs instead:
$\left|V_{(y,d)=g}\right|=h\left|S_{(y,d)=g}\right|,\ \ \forall g\in\mathcal{Y\times D}.$ (2)
For the remainder of this section, the latter set of constraints (2) are assumed, although (1) can easily be substituted if desired (as this would further increase the MMD between $T_{X}$ and $V_{X}$), or if domain labels are unavailable.
It is also proposed that $V$ should be maximally domain shifted with respect to $T$. It is proposed to measure this as the MMD between the distributions of feature sets $\theta_{F}\left[T_{X}\right]$ and $\theta_{F}\left[V_{X}\right]$, where $\theta_{F}[\cdot]$ denotes the image under $\theta_{F}$.
@@
$\mu_{\mathbb{P}_{T}}=\frac{1}{\left|T_{X}\right|}\sum_{f\in\theta_{F}\left[T_{X}\right]}{\phi(f)}.$
$\text{MMD}\left(\mathbb{P}_{T},\mathbb{P}_{V}\right)=\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}.$
1.3 Problem formulation
The partitioning problem can now be formulated as
$\arg{\max_{T,V}\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}}$ (3)
subject to (2).
This is equivalent to minimising the standard kernel k-means objective, subject to the same constraint:
$\arg{\min_{T,V}{\sum_{\begin{subarray}{c}J\in\left\{T,V\right\}\\ f\in\theta_{F}\left[J_{X}\right]\end{subarray}}\left\|\phi(f)-\mu_{\mathbb{P}_{J}}\right\|_{\mathcal{H}}^{2}}}$ (4)
subject to (2).
Proof.
The relation can be derived by applying the law of total variance. The objective in (4) is to minimise the within-cluster sum of squared deviations from each centroid (in feature space). As the total variance is constant, this is equal to maximising the between-cluster sum of squares from each cluster centroid to the grand one, weighted by cluster size:
$\arg{\max_{T,V}{\sum_{J\in\{T,V\}}{\left|J_{X}\right|\left\|\mu_{\mathbb{P}_{J}}-\mu_{\mathbb{P}_{S}}\right\|_{\mathcal{H}}^{2}}}}.$
Since the cluster sizes are fixed by the constraint, the $\left|J_{X}\right|$ term can be dropped from the objective function. With only two centroids, this is then equivalent to directly maximising the distance between them, which is the formulation in (3). ∎
(4) can be solved by iterating between 2 steps until convergence:
@@ 1.3 Problem formulation
    The update (maximisation) step computes the distance matrix $D\in\mathbb{R}^{n\times 2}$ from each point to each centroid using the kernel trick. For large datasets, the Nyström method [Chitta2014ScalableK-means] is employed to reduce the complexity of the kernel computations from $O(n^{2})$ to $O(qn)$, by computing only a randomly selected submatrix of the full kernel, of size $q\times n$.
@@
    The constrained assignment (expectation) step computes the one-hot cluster assignment matrix $U\in\left\{0,1\right\}^{n\times 2}$ that assigns each point to exactly one of the two centroids.
As in the standard Lloyd’s algorithm, convergence to a local minimum is guaranteed as each step individually optimises (4). $U$ is the solution to the binary linear program (LP):
$\mathrm{arg}\,\min_{U}$ $\,{\operatorname{Tr}\left(U^{T}D\right)}$ (5)
@@
$U_{ij}\in\left\{0,1\right\}$ $\operatorname{for}i\in\{1,\ldots,n\};\ j\in\left\{1,2\right\}$ (6)
$\sum_{j}U_{ij}=1$ $\operatorname{for}i\in\{1,\ldots,n\}$ (7)
$\sum_{i|\left(y_{i},d_{i}\right)=g}U_{ij}=\operatorname{round}\left(h\left|S_{(y,d)=g}\right|\right)$ $\operatorname{for}g\in\mathcal{Y\times D};\ j=1\ \text{or}\ 2.$ (8)
@@
The constraints satisfy Hoffman’s sufficient conditions for total unimodularity [Heller1956AnDantzigs] (in particular, it can be seen that (7) and (8) form two disjoint sets of constraints, and every element of $U$ is referenced at most once in each set). The consequence is that the LP will always have integer solutions, without having to enforce them explicitly. This means the binary constraint (6) can be relaxed to
$0\leq U_{ij}\leq 1\qquad\operatorname{for}\ i\in\{1,\ldots,n\};\ j\in\{1,2\}$
and the problem can be solved without integer constraints.
To enforce soft constraints, (8) can be replaced by the inequality
$\operatorname{round}\left(h\left(1-\tau_{g}\right)\left|S_{(y,d)=g}\right|\right)\ \leq\sum_{i|\left(y_{i},d_{i}\right)=g}U_{ij}\leq\operatorname{round}\left(h\left(1+\tau_{g}\right)\left|S_{(y,d)=g}\right|\right)\ \\ \operatorname{for}\ g\in\mathcal{Y\times D;\ }j=1\ \text{or}\ 2,$

### v3

# Clustering-Based Validation Splits for Model Selection under Domain Shift

@@ Abstract
This paper considers the problem of model selection under domain shift. Motivated by principles from distributionally robust optimisation and domain adaptation theory, it is proposed that the training-validation split should maximise the distribution mismatch between the two sets. By adopting the maximum mean discrepancy (MMD) as the measure of mismatch, it is shown that the partitioning problem reduces to kernel k-means clustering. A constrained clustering algorithm, which leverages linear programming to control the size, label, and (optionally) group distributions of the splits, is presented. The algorithm does not require additional metadata, and comes with convergence guarantees. In experiments, the technique consistently outperforms alternative splitting strategies across a range of datasets and training algorithms, for both domain generalisation and unsupervised domain adaptation tasks. Analysis also shows the MMD between the training and validation sets to be well-correlated with test domain accuracy, further substantiating the validity of this approach.
@@ 1 Introduction
The ability for models to maintain high performance on data lying outside their training distribution, known as domain generalisation (DG), is crucial to the widespread deployment of artificial intelligence. Although extensive research has been conducted towards developing more generalisable training algorithms (Gulrajani & Lopez-Paz, 2021), significantly less focus has been given to increasing the robustness of the model selection process, despite being as integral a part of the learning problem, and indeed, just as susceptible to distribution shifts, as the fitting of the model itself.
As with model parameters, hyperparameters chosen based on in-distribution (ID) performance lack optimality guarantees on out-of-distribution (OOD) test data. Metadata-based dataset splitting, which creates OOD validation sets distinct from both the training and test data, is commonly used in this scenario, and has empirically been shown to encourage the selection of more generalisable hyperparameters (Koh et al., 2021). This paper is motivated by the principles of distributionally robust optimisation (DRO) (Rahimian & Mehrotra, 2019), which aims to optimise for worst-case performance within some uncertainty set of distributions. This suggests that the validation set should indeed be maximally domain shifted from the training set, while still retaining a relevant distribution. It is proposed that a worst-case dataset split which maximises domain mismatch would balance these two aims, and further encourage the selection of robust hyperparameters.
@@
Among other considerations, the measure of mismatch should be such that the partitioning problem can be tractably solved. To this end, it is noted that performing kernel k-means clustering is equivalent to maximising the empirical maximum mean discrepancy (MMD) between clusters (weighted by cluster size). Thus, this paper proposes to perform a validation split based on kernel k-means, and presents a modified clustering algorithm for this purpose. Specifically, constraints are introduced to the cluster assignment step to control the holdout fraction (i.e., the cluster sizes), and to preserve class and (optionally) domain/group distributions; this is then formulated and solved as a linear program, providing convergence guarantees not present in prior work. In addition, the Nyström method for low-rank approximation is employed to make the algorithm tractable for large datasets.
@@
The proposed method provides a model selection strategy based on OOD performance that does not require additional metadata, which is not always available. It is also argued that the method is able to capture more nuances in real-world data than can be described by a single domain variable, which is another limitation of metadata-based splits. For example, in tumour identification, domain shifts can be caused by variations in sample preparation methods or patient populations. Or, in wildlife monitoring (both acoustic and visual), these can be due to differences in environmental or weather conditions, data collection equipment, or new unseen events (Koh et al., 2021). However, all this information is unlikely to be described in the available metadata, which may only state the hospital of origin or the recording location, respectively. In such cases, splitting the data along lines closer to the underlying cause of the shift, as determined by the clustering algorithm, rather than a more loosely correlated proxy, may permit a more informed model selection which can result in hyperparameters better suited to the nature of the shift.
This paper considers the two most common paradigms for learning under domain shift: domain generalisation (DG) and unsupervised domain adaptation (UDA). Both paradigms assume the data can be grouped into “domains” which share certain characteristics, and that each domain is associated with a different data distribution. Given multiple source domains, DG seeks to learn domain-invariant feature representations such that the model becomes robust to future (unknown) changes in the data. In contrast, UDA assumes unlabelled data from a specific target domain is also available during training. Although DG operates in a stricter setting, it also requires fewer assumptions about the target data – for example, UDA methods often assume minimal label shift, but this can have a significant detrimental effect on performance (Napoli & White, 2024). UDA is also less stable than DG, as it is susceptible to phenomena such as catastrophic forgetting, negative transfer, or overfitting on the target data. However, DG requires that the causal mechanisms behind the domain shift remain stable, that is, that the test domain is drawn from the same meta-distribution as the source domains. As this is difficult to ascertain in practice, test domain performance is also harder to predict. In contrast, test domain accuracy in UDA is directly coupled to domain alignment quality, given the theoretical link between the two (Redko et al., 2022).
@@
  • Description of a constrained clustering algorithm based on kernel k-means which can be used to perform a training-validation split in applications expected to involve domain shifts.
  • Comparison of this approach with existing validation splitting strategies for a range of datasets and algorithms, in both DG and UDA settings.
@@ 2 Prior work
Gulrajani & Lopez-Paz (2021) reviewed 3 criteria for model selection in a DG setting: ID accuracy (on a randomly held-out subset of each training domain); OOD accuracy on an additional domain held-out using metadata (referred to as leave-one-domain-out); and test-distribution accuracy on a held-out subset of the test domain (referred to as the oracle criterion), which can be used to provide an upper bound on performance. Where multiple validation domains are given, a DRO-style treatment of model selection which considers only the worst-performing domain is studied by Sagawa et al. (2019); Gao et al. (2023); Pfohl et al. (2022) – although this again relies on the availability of metadata.
@@
In addition to validation accuracy, it has been suggested that a model’s stability to distribution shifts should also be explicitly considered. Prior work has quantified stability in terms of the expected calibration error (the average deviation between accuracy and confidence) (Wald et al., 2021); the MMD between features from different domains (Lyu et al., 2023); or the average variation of each feature between domains (Ye et al., 2021).
It has also been proposed to induce a domain shift between the training and validation sets by performing mix-up augmentation on the held-out data (Lu et al., 2023). This links to a range of general robustness benchmarks, where the evaluation sets have been subjected to various synthetic transformations. For example, visual corruptions and perturbations (Hendrycks & Dietterich, 2019), stylisation (Geirhos et al., 2018), the addition of spurious cues (Li et al., 2022), and adversarial filtration (Hendrycks et al., 2019) have all been employed; a more complete review of this approach is given in Koh et al. (2021).
Non-random data splits have previously been used to produce domain-shifted evaluation sets, but no prior work has investigated these in a model selection context. Søgaard et al. (2021) proposed an adversarial splitting heuristic based on k-nearest neighbours and the Wasserstein distance between term count distributions of text documents, but this is not generally applicable to features in $\mathbb{R}^{n}$, and no attempt was made to control the label distribution. Adversarial data splits have also been used in meta-learning to improve DG (Gu et al., 2023; Wang et al., 2024). Recently, Napoli & White (2025) proposed a data selection algorithm which can be used to choose a validation set closer to the test domain in the UDA setting. However, this is not applicable if the split needs to be decided before the test domain is accessible, e.g., for test-time adaptation.
Finally, with an approach most similar to ours, Wecker et al. (2020) proposed an evaluation split based on k-means clustering, with constraints to control for cluster size and label distribution. Our paper develops on this work in a number of ways. Firstly, the motivation of Wecker et al. (2020) is merely to create more challenging data splits, and is not applied to improving model selection nor theoretically connected to distributional robustness. Secondly, Wecker et al. (2020) impose the constraints using a greedy strategy, which are susceptible to local minima. In contrast, we solve a constrained assignment problem using linear programming, which avoids this effect and results in better quality solutions. We also generalise the label-wise constraints to domains for multi-domain data. Finally, our theoretical analysis motivates a kernelised version of this algorithm, thereby exploiting the relation between the kernelised objective (which we link to the MMD) and validation accuracy.
The theoretical relation between the MMD and kernel k-means clustering has also been derived in other contexts. For example, França et al. (2017) obtain similar results from the viewpoint of generalised energy statistics, while Ohl et al. (2024) use the relation to improve training of unsupervised decision trees.
3 Preliminaries and notation
@@
Let $S=\{(x_{1},y_{1},d_{1}),\ldots,(x_{n},y_{n},d_{n})\}$ be the development set consisting of input-label-domain triplets over $\mathcal{X}\times\mathcal{Y}\times\mathcal{D}^{S}$. Similarly, let $E$ be the evaluation set over $\mathcal{X}\times\mathcal{Y}\times\mathcal{D}^{E}$, where $\mathcal{D}^{S}\cap\mathcal{D}^{E}=\emptyset$ (i.e., the development and evaluation domains are disjoint), and let $\mathcal{Y}$, $\mathcal{D}^{S}$ and $\mathcal{D}^{E}$ be finite sets. For ease of notation, subscripts are used on sets to simplify set-builder notation, in two ways. Firstly, “slices” of a set are denoted using capitals, for example:
@@
$\theta^{*}=\arg{\min_{\theta\in\Theta}R(\theta,V)},$ (1)
where $R(\theta,V)$ is the error of $\theta$ on $V$, and $\theta^{*}$ is evaluated on $E$.
3.1 Theoretical motivation
DRO (Rahimian & Mehrotra, 2019; Sagawa et al., 2019) is a well-established and theoretically grounded paradigm for obtaining more generalisable solutions to optimisation problems. Instead of reducing overall error, DRO adopts a minimax strategy which minimises the worst-case error over some uncertainty set $\mathcal{Q}$. By defining $\mathcal{Q}$ as the possible (suitable) partitions of $S$, the model selection problem becomes
$\min_{\theta\in\Theta}\max_{\{T,V\}\in\mathcal{Q}}R(\theta,V).$ (2)
This could be solved directly via an expensive alternating optimisation process (Gu et al., 2023). However, we propose instead to find $\arg\max_{T,V}R(\theta,V)$, the worst-case partition in the sense of $R$, in a single step by solving a surrogate problem based on the well-known relationship between $R(\theta,V)$ and the mismatch between $T$ and $V$, which has been fundamental to domain adaptation theory (Ben-David et al., 2006; 2010). By using the MMD to measure this mismatch, the partitioning problem can be efficiently solved through clustering, as will be shown in Section 4. The theoretical relation between $\text{MMD}(T,V)$ and $R(\theta,V)$ is given by the generalisation error bound (Redko et al., 2022, Theorem 36)
$R(\theta,V)\leq R(\theta,T)+\text{MMD}(T,V)+\lambda$ (3)
where $\lambda$ is a term depending on the capacity of the hypothesis space and the combined performance of an ideal model on both $T$ and $V$. A linear correlation has also been observed empirically (Napoli & White, 2025). Although maximising the MMD is not guaranteed to maximise this bound due to the dependence of $R(\theta,T)$ on the partitioning, the empirical analysis in Section 5.5 shows that this term remains fairly stable in practice.
Equation (3) is quite general and not strictly limited to covariate shift. However, the robustness of the bound relies on the ability of the model to simultaneously perform well on both $T$ and $V$ (through $\lambda$), and this can be affected by the presence of conditional shifts. There are also assumptions made about both the (joint) data distributions and the hypothesis class of models – we refer to Redko et al. (2022) for formal statements of these – and these ultimately define the reproducing kernel Hilbert space in which the MMD in the bound is measured. Thus, by symmetry, the choice of kernel shapes the underlying assumptions about the model and data, and ultimately determines whether the MMD being optimised is predictive of the discrepancy between the training and validation losses.
4 Method
This section defines the constraints needed to ensure an appropriate split, formulates the partitioning problem, and describes an algorithm to solve the optimisation.
$T$ and $V$ should be of sizes determined by a user-defined holdout fraction $h$ satisfying $0<h<1$, and have equivalent class distributions. This can be achieved by constraining the size of each label group in $V$ to $h$ times the size of the corresponding group in $S$:
This constraint ensures that there are sufficient examples from each class in both $T$ and $V$ to properly train and validate the models. Without it, the data may have a tendency to cluster by class, which may result in certain classes being placed entirely in one set.
It may also be necessary or desirable to control domain distributions. For example, certain training algorithms may require that the domains in $T$ be uniformly represented to avoid overfitting; controlling the distributions of validation groups has also been suggested to reduce noise in the hyperparameter tuning process (Sagawa et al., 2019). Finally, in the multi-task learning setting (i.e., some or all of the domains correspond to different tasks, with separate validation metrics), the domain split-ratio should also be controlled: this both prevents bias towards a specific task, and also ensures that each task has sufficient representation in the validation set to reliably estimate the corresponding validation metric. In these cases, the constraints should be taken over all $(y,d)$ pairs instead:
$\left|V_{(Y,D)=g}\right|=h\left|S_{(Y,D)=g}\right|,\ \ \forall g\in\mathcal{Y}\times\mathcal{D}^{S}.$ (5)
For the remainder of this section, the latter set of constraints (5) are assumed, although (4) can easily be substituted if desired (as relaxing the constraints will increase the clustering objective), or if domain labels are unavailable.
The objective is to maximise the discrepancy between $T$ and $V$, measured using the MMD. Given the potentially high dimensionality of $\mathcal{X}$, doing this directly in the input space can be impractical. Therefore, it is proposed to instead measure the MMD between the distributions of feature sets $\theta_{F}\left[T_{X}\right]$ and $\theta_{F}\left[V_{X}\right]$, where $\theta_{F}[\cdot]$ denotes the image under $\theta_{F}$. Moreover, these intermediate features are those most highly correlated with the model’s outputs, meaning this is the representation space for which the MMD is most predictive of classification error. For further intuition, consider that a large MMD in this space implies a domain shift that the model is not invariant to (and thus is more likely to cause problems) – hence why this is also the representation space normally targeted by feature alignment methods (Gulrajani & Lopez-Paz, 2021). For models where no intermediate representations are available, the input representations can be used directly, or an alternative feature extractor or dimensionality reduction technique can be used. Prior knowledge on the expected nature of the domain shift could also be incorporated into the clustering features; a detailed discussion this is given in Appendix A.
@@ 4 Method
$\mu_{\mathbb{P}_{T}}=\frac{1}{\left|T_{X}\right|}\sum_{f\in\theta_{F}\left[T_{X}\right]}{\phi(f)}.$ (6)
$\text{MMD}\left(\mathbb{P}_{T},\mathbb{P}_{V}\right)=\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}.$ (7)
The partitioning problem can now be formulated as
@@
This is equivalent to performing kernel k-means clustering with $k=2$ (i.e., using one cluster for each of $T$ and $V$), subject to the same constraints:
[…]
where $\Psi(T,V)=\operatorname{SSq}(T_{X})+\operatorname{SSq}(V_{X})$ is the standard kernel k-means objective function and
$\operatorname{SSq}(T_{X})=\sum_{f\in\theta_{F}\left[T_{X}\right]}\left\|\phi(f)-\mu_{\mathbb{P}_{T}}\right\|_{\mathcal{H}}^{2}$ (10)
is the sum of squared deviations of a set of points from their centroid (in feature space).
Theorem 1.
Problems (8) and (9) are equivalent.
Proof sketch.
The equivalence can be derived by applying an ANOVA sum-of-squares decomposition, followed by a substitution based on the polarisation identity. The resulting identity is
$\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}=\frac{|S_{X}|}{|T_{X}||V_{X}|}(\operatorname{SSq}(S_{X})-\Psi(T,V)).$ (11)
For a full derivation, see Appendix B. As $S_{X}$ is constant with respect to the cluster allocations, and the cluster sizes are fixed by the constraint, these terms can all be dropped from the objective function. Hence, (8) and (9) are equivalent. ∎
Problem (9) can be solved by applying a variation of Lloyd’s algorithm (Chitta et al., 2014).
Algorithm 1
Constrained kernel k-means clustering
Given an initial set of assignments, alternate between 2 steps until convergence or max iterations reached:
@@
    Distance Update (maximisation). Compute the distance matrix $D\in\mathbb{R}^{n\times 2}$ from each point to the centroid of each cluster using the kernel trick. For large datasets, use the Nyström method (Chitta et al., 2014) to reduce the complexity of the kernel computations from $O(n^{2})$ to $O(qn)$, by computing only a randomly selected submatrix of the full kernel, of size $q\times n$.
@@
    Constrained Assignment (expectation). Compute the one-hot cluster assignment matrix $U\in\left\{0,1\right\}^{n\times 2}$ that assigns each point to exactly one of the two clusters. $U$ is the solution to the binary linear program (LP):
    $\arg\min_{U}{\sum_{i,j}U_{ij}D_{ij}}$ (12)
@@
    $U_{ij}\in\left\{0,1\right\}\ \quad$ $\operatorname{for}\ i\in\{1,\ldots,n\};\ j\in\left\{1,2\right\}$ (13)
    $\sum_{j}U_{ij}=1\ \quad$ $\operatorname{for}\ i\in\{1,\ldots,n\}$ (14)
    $\sum_{i|\left(Y_{i},D_{i}\right)=g}U_{ij}=\operatorname{round}\left(h\left|S_{(Y,D)=g}\right|\right)\ \quad$ $\operatorname{for}\ g\in\mathcal{Y}\times\mathcal{D}^{S};\ j=1\ \text{or}\ 2.$ (15)
@@
The constraints satisfy Hoffman’s sufficient conditions for total unimodularity (Heller & Tompkins, 1956) (in particular, it can be seen that (14) and (15) form two disjoint sets of constraints, and every element of $U$ is referenced at most once in each set). The consequence is that the LP will always have integer solutions, without having to enforce them explicitly. This means the binary constraint (13) can be relaxed to
$0\leq U_{ij}\leq 1\operatorname{for}\ i\in\{1,\ldots,n\};\ j\in\{1,2\}$ (16)
and the problem can be solved without integer constraints. To enforce soft constraints, (15) can be replaced by the inequality
$\operatorname{round}\left(h\left(1-\tau_{g}\right)\left|S_{(Y,D)=g}\right|\right)\ \leq\sum_{i|\left(Y_{i},D_{i}\right)=g}U_{ij}\leq\operatorname{round}\left(h\left(1+\tau_{g}\right)\left|S_{(Y,D)=g}\right|\right)\ \operatorname{for}\ g\in\mathcal{Y}\times\mathcal{D}^{S};\ j=1\ \text{or}\ 2,$ (17)
@@
Proposition 2.
Algorithm 1 converges to a locally optimal partitioning in a finite number of iterations.
Proof sketch.
Note that $\Psi(T,V)$ is bounded below by 0 and is also non-increasing, since both the centroid updates and cluster assignments are (or can be interpreted as) optimisation problems which share the same objective function as (9), can be solved globally at each iteration, and do not violate any of the constraints. Note also that only a finite number of partitionings are possible, meaning $\Psi(T,V)$ can only decrease a finite number of times. Therefore, convergence in finite time is guaranteed. For a full proof, see Bradley et al. (2000). ∎
Although Proposition 2 provides a trivial upper bound for the number of iterations required until convergence, Lloyd’s algorithm is known to be fast in practice, and the runtime is often considered to be linear in the number of datapoints based on empirical observations (Cordeiro de Amorim & Makarenkov, 2023). For further theoretical analysis on the number of iterations required, see, e.g., Arthur et al. (2009); Arthur & Vassilvitskii (2006).
4.1 Kernel choice
As noted in Section 3.1, the choice of kernel implies certain assumptions about the data. In some applications (e.g., physical modelling), known feature constraints naturally guide the kernel choice. In our case, however, only relatively weak geometric assumptions can be made. The kernel should also be characteristic to ensure the MMD can distinguish between any two distributions (Sriperumbudur et al., 2009). However, we note that, unlike kernel methods which make use of the kernel distances directly, kernel k-means clustering only compares between kernel distances when assigning datapoints, which we would argue makes the solution less sensitive to kernel choice or kernel hyperparameters. Additionally, the cluster size constraints restricts the solution space, which further desensitises the solution to kernel choice.
For the experiments in the subsequent section, we opt to use a Gaussian radial basis function (RBF) kernel $\kappa(x,y)=e^{-\gamma\left\|x-y\right\|^{2}}$, which is a convenient and popular choice for MMD estimation (Sriperumbudur et al., 2009). Perhaps more relevantly, we also note that the correlation between the MMD (with an RBF kernel) and $R(\theta,V)$ has been supported by empirical observations (Napoli & White, 2025).